When 2 resistances r1 and r2 are connected in parallel to a source of emf the ratio of heat produced is

Grade:12

1 Answers

Saswat Pani
6 years ago
As the circuit is connected in parallelIn R1 I1 = V/R1In R2 I2 = V/R2So H1/H2 = (V^2/R1^2 × R1×t)/ (V^2/R2^2 × R2×t)=R2/R1